I need to use substitution method
15x-10y=1
5y=-2+15x

Found 3 solutions by mananth, ikleyn, greenestamps:
Answer by mananth(16949)

15 x + -10 y = 1 .............1

5 y = -2 15 x ---------2
/ 5 =
y = -0.4 3 x
Plug the value of y in (1)
15x-10(-0.4+3x)
15x+4-30x=1
15x-30x=1-4
-15x=-3
/-15
x=0.2
Plug the value of x in (1)
15x-10y=1
15*0.2-10y=1
3-10y=1
-10y=-2
/-10
y=0.2

Answer by ikleyn(53749)

.
I need to use substitution method
15x - 10y = 1
5y = -2 + 15x
~~~~~~~~~~~~~~~~~~~~~~~~


        It can be solved in much more efficient way than @mananth does it in his post.

        I will apply a  " smart "  substitution method.


Your starting equations are 

    15x - 10y = 1    (1)
    5y = -2 + 15x    (2)


From equation (2), express  15x = 5y + 2 and substitute it as a whole block into equation (1), 
replacing the term 15x there


    (5y + 2) - 10y = 1,

     5y - 10y = 1 - 2,

        -5y   =   -1

          y   =   1/5 = 0.2.


Now substitute y = 0.2 into equation (1) and find x


    15x - 10*0.2 = 1,

    15x - 2 = 1,

    15x = 1 + 2 = 3,

      x = 3/15 = 0.2.


ANSWER.  x = 0.2,  y = 0.2.

Solved.

As you see,  in this case  " smart "  substitution works more effectively
than a standard  " stupid "  substitution.


The meaning of this problem for you is to learn advantages of "smart" substitution method.



Answer by greenestamps(13327)



Pay close attention to the solution from tutor @ikleyn.

In a typical algebra class, given a system of two linear equations, you are taught to solve one of the equations for one of the variables (by itself) and substitute the result in the other equation.

But as the tutor shows in this problem, the solution is much easier because the expression "15x" appears in both of the given equations; so solving one of the equations for "15x" and substituting in the other equation makes the solution easy.